# for Verifiable Chain-of-Thought Financial Reasoning

### v2

@@ 6 Conclusions and Future Work
We acknowledge several limitations in this work that we plan to address in future research.
First, our dataset is entirely synthetic, generated from symbolic templates. While this design enables fine-grained control, contamination-free generation, and automatic verifiability, it may lack the linguistic diversity and contextual richness of real-world financial texts. In future work, we plan to incorporate real-world financial documents, such as earnings reports, investor communications, or financial news, as seed sources for semi-structured or templated generation, allowing us to better simulate naturally occurring language while preserving symbolic grounding.
Second, the benchmark focuses narrowly on symbolic numerical reasoning and does not capture qualitative, contextual, or strategic aspects of financial decision-making, such as risk assessment, or market sentiment. To address this, we aim to complement our benchmark with additional tasks that evaluate models on these higher-level reasoning dimensions, potentially through multi-modal inputs (e.g., combining text with charts or scenarios) or interactive decision-making simulations.
Third, FinChain is restricted to English and U.S. centric financial conventions, such as currency formats, and investment norms, which limits its generalizability to multilingual and regional financial contexts. In future extensions, we plan to expand coverage to other languages and financial systems, through collaboration with domain experts and native speakers to localize templates and ensure culturally grounded reasoning tasks.
Fourth, we use regular expressions for parsing responses of LLMs, but it is not ideal. That is why ChainEval may be computed on unrelated parts of the text. In future we plan to migrate to LLM-based parsing of responses and inclusion of length penalty, which will help to penalize models for overly verbose steps.
@@ Acknowledgments
We thank all our financial experts (Salim Tlemçani, Alfred Choi, Petrus Kung, Shaobo Wang, Bowen Hao, Ruofei Li, and Xunwen Zheng) for their assistance in reviewing the templates. This project would not have been possible without their support. We also express our sincere gratitude to Muhammad Usman Safder and Ayesha Gull for their interest in this project and for developing the live demo and demonstration website. Finally, we thank the reviewers for their valuable feedback, which significantly improved the quality of this paper.
Appendix A Financial-Symbolic: Template Creation Prompt
@@
Appendix B Financial Expert Demography
@@ Appendix C Annotation Platform
Appendix D Annotated Template Issue Statistics
@@
Appendix E Model Detail Information
We provide the details of the evaluated models in Table 3.
Appendix F Review Rubrics
@@
F.1 Reasoning Process Quality
@@
  • 2 (Poor): Some reasoning attempt but with major factual or procedural flaws; inconsistent or unclear chain of thought.
@@
F.2 Final Answer Accuracy
@@
Appendix G Metric Evaluation and Ablations
To better understand the behavior of our proposed metric, we conducted a series of ablation experiments and comparative analyses. All quantitative results reported in this section are benchmarked against expert human evaluations of reasoning quality (see Appendix B for expert details).
G.1 Ablations of the DTW-Based Metric
@@
G.2 Comparative Evaluation
We also evaluated a range of traditional text-similarity and reasoning metrics, including ROUGE-2, ROUGE-L, step-level precision and recall (marked as ‘non-DTW’ in the table). Each metric was correlated with expert-assigned Reasoning Process Quality score. Table 5 summarizes the top Spearman correlations with expert process judgments.
@@ G.2 Comparative Evaluation
Table 5: Spearman correlation with expert process scores across all metrics.
G.3 Discussion
The DTW-based variants consistently achieve the highest correlation with expert judgments, with the Normalized DTW Alignment Score (Gate Mode) emerging as the most reliable indicator of reasoning faithfulness. The “Soft” variant yields slightly lower but still strong correlations, suggesting that the gating formulation better captures strict consistency, while the soft variant provides smoother sensitivity to near-correct reasoning. Compared to traditional metrics such as ROUGE or simple step-level precision and recall, DTW captures not only semantic similarity but also the structural coherence and numerical consistency of reasoning chains. These results highlight the usefulness of our proposed metric.
G.4 Error Analysis
@@
We observed that longer and more verbose solution steps tend to receive higher scores, primarily because they exhibit greater similarity to our gold-standard solutions, likely due to redundancy. Additionally, some models (such as Qwen 2.5 Instruct) often produce repetitive steps, which reduces the accuracy of our current regular-expression-based step parsing.
To address these issues, future iterations of our metric will incorporate LLM-based answer parsing and introduce a length penalty to better account for verbosity and repetition.

### v3

@@ 6 Conclusion and Future Work
This work has several limitations, which can be addressed in future research. First, our dataset is entirely synthetic and generated from symbolic templates. While this design enables controllable, contamination-free generation and automatic verification, it may lack the linguistic diversity and contextual richness of real-world financial texts. Future work could incorporate real financial documents as seed inputs for semi-structured generation while preserving symbolic grounding.
Second, the benchmark focuses on symbolic numerical reasoning and does not capture qualitative or strategic aspects of financial decision-making, such as risk assessment or market sentiment. Extending the benchmark to cover these higher-level reasoning dimensions remains an open challenge.
Third, FinChain is limited to English and primarily reflects U.S.-centric financial conventions, which restricts its applicability to multilingual and regional contexts. Expanding to additional languages and financial systems is an important direction for future work.
Finally, our evaluation relies on automatic parsing of model-generated reasoning chains, which can be sensitive to formatting variations or extraneous text. Improving the robustness of step-level alignment through more structured output formats or tighter integration with symbolic execution is another possible direction for future work.
@@ Acknowledgments
We thank all our financial experts (Salim Tlemçani, Alfred Choi, Petrus Kung, Shaobo Wang, Bowen Hao, and Xunwen Zheng) for their assistance in reviewing the templates. This project would not have been possible without their support. We also express our sincere gratitude to our student volunteers Muhammad Usman Safder and Ayesha Gull for their interest in this project and for developing the live demo and demonstration website. Finally, we thank the reviewers for their valuable feedback, which helped us improve the quality of this paper.
Appendix A FinChain Construction Details
A.1 Template Creation Prompt
@@
A.2 Template Examples
Here, we present example templates of three compound interest (CI) financial questions, grouped by difficulty level, including basic, intermediate, and advanced.
Basic Level.
def template_ci_quarterly_compounding():
"""Basic: Compound Interest with Quarterly Compounding"""
investor_name = random.choice(investor_names)
project_name = random.choice(project_names)
# Parameters
principal = random.randint(1_000, 7_000) # \$
rate = round(random.uniform(2, 8), 2) # annual %, two decimals
time = random.randint(1, 3) # years
n = 4 # quarterly
# ---------- Question ----------
question = (
f"{investor_name} invests \${principal} in {project_name}. "
f"The account earns {rate:.2f}% interest per year, compounded quarterly, "
f"for {time} years. What is the total compound interest earned "
# ---------- Reasoning ----------
# Step 1: future (compound) amount
future_value = principal (1 + rate / (100 n)) (n time)
# Step 2: compound interest
ci = future_value - principal
# Round only for display
fv_display = f"\${future_value:.2f}"
ci_display = f"\${ci:.2f}"
# ---------- Solution ----------
solution = (
[…]
f" = {ci_display}"
return question, solution
Intermediate Level.
def template_ci_rate_and_total_known():
"""Intermediate: Compound Interest with nominal rate, time, and frequency known"""
investor_name = random.choice(investor_names)
project_name = random.choice(project_names)
# ---------- Parameters ----------
total_amount = random.randint(5_000, 15_000) # Final amount A (\$)
rate = round(random.uniform(2, 10), 2) # Nominal annual rate %
time = random.randint(1, 5) # Years
freq_name, n = random.choice(
[("semi-annually", 2), ("quarterly", 4), ("monthly", 12)]
# ---------- Question ----------
question = (
f"{investor_name} received a total amount of \${total_amount:,.2f} "
f"from their investment in {project_name}. "
f"The investment grew at a nominal annual interest rate of {rate:.2f}% "
f"compounded {freq_name} for {time} years. "
f"Calculate the compound interest earned (in dollars)."
# ---------- Reasoning ----------
# Step 1: periodic rate and growth factor
periodic_rate = round(rate / 100 / n, 6) # r_p
growth_factor = round((1 + periodic_rate) (n time), 6)
# Step 2: principal P
principal = round(total_amount / growth_factor, 2) # 2 dp dollars
# Step 3: compound interest CI
ci = round(total_amount - principal, 2)
# ---------- Solution ----------
solution = (
[…]
f" P = A $\div$ growth factor = "
[…]
f" CI = A - P = \${total_amount:,.2f} - \${principal:,.2f} = \${ci:,.2f}"
return question, solution
Advanced Level.
def template_ci_with_additional_deposit():
"""Advanced: Compound Interest with a Mid Term Additional Deposit (needs 4 steps)"""
investor_name = random.choice(investor_names)
project_name = random.choice(project_names)
# --- parameters ---
principal = random.randint(2000, 8000) # initial \$
rate = round(random.uniform(3, 10), 2) # % p.a.
time = random.randint(3, 7) # total years (>=3 so a mid deposit makes sense)
n = random.choice([1, 2, 4, 12]) # compounds per year
deposit = random.randint(500, 4000) # extra \$
deposit_time = random.randint(1, time - 1) # year when deposit is made
# ---------- Question ----------
question = (
f"{investor_name} initially invested \${principal} in {project_name} at an annual "
f"rate of {rate:.2f}%, compounded {n} times a year, for a total of {time} years. "
f"Exactly {deposit_time} years after the start, they added an extra \${deposit} "
f"to the same account under the same rate and compounding frequency. "
f"Calculate the total compound interest earned by the end of the {time} years."
# ---------- Reasoning ----------
# Step 1 - periodic rate
periodic_rate = round(rate / (100 n), 4)
# Step 2 - grow the original principal for the full period
periods_principal = n time
fv_principal = round(principal (1 + periodic_rate) periods_principal, 2)
# Step 3 - grow the later deposit for the remaining (time - deposit_time) years
remaining_years = time - deposit_time
periods_deposit = n remaining_years
fv_deposit = round(deposit (1 + periodic_rate) periods_deposit, 2)
# Step 4 - combine amounts and find compound interest
total_future_value = round(fv_principal + fv_deposit, 2)
total_contributions = principal + deposit
compound_interest = round(total_future_value - total_contributions, 2)
# ---------- Solution ----------
solution = (
[…]
f" Total FV = FV1 + FV2 = \${fv_principal:.2f} + \${fv_deposit:.2f} = "
[…]
f" Compound Interest = Total FV - Contributions = "
f"\${total_future_value:.2f} - \${total_contributions} = \${compound_interest:.2f}"
return question, solution
A.3 Data Validation Criteria
Directly prompting large language models to generate symbolic financial reasoning templates can lead to inconsistencies or incomplete specifications. To address these issues, we apply the following validation constraints prior to expert review.
Cross-national inconsistencies.
Generated questions occasionally contained country-specific financial conventions (e.g., currencies, indices, or terminology). All such cases are standardized to U.S.-based financial settings.
Precision mismatch.
In some cases, displayed values were rounded while computations used full precision. We align computational outputs with the displayed numerical precision.
Incomplete input specification.
Some questions omitted variables required for computation. These cases are revised to include all necessary inputs.
Unit consistency.
Currency symbols and units were inconsistently applied across questions and solutions. All templates are standardized to consistent units.
Non-informative steps.
Certain generated solutions decomposed simple calculations into trivial substeps or omitted intermediate reasoning. These solutions are revised to reflect substantive reasoning steps.
Multiple targets.
Some templates requested multiple output values, complicating evaluation. We constrain all templates to require a single target value.
A.4 Expert Review and Annotation Protocol
To further enhance data quality, we recruited ten financial experts to review all validated templates. The expert panel consists of seven graduate students in economics, finance, and related quantitative disciplines, and three industry professionals with experience in quantitative research, financial engineering, and risk management. Annotators were selected through an internal vetting process to ensure domain expertise and professional credibility. Additional demographic details are provided in § B.1.
Annotation Platform.
We developed a Streamlit-based annotation platform to facilitate efficient expert review. Implementation details are provided in § B.2.
Pilot Study.
Prior to full annotation, we conduct a pilot study in which 20 templates are reviewed by all annotators to calibrate evaluation standards. After calibration, all annotators agreed on the correctness of the pilot templates.
Main Annotation.
The remaining 270 templates are randomly distributed among annotators, with each template reviewed by a single expert. Out of 290 total templates, 29 are identified as incorrect and subsequently revised by financial experts. Summary statistics of identified issues are reported in § B.3.
Appendix B Annotation and Quality Control
B.1 Financial Expert Demography
@@ B.2 Annotation Platform
B.3 Annotated Template Issue Statistics
@@
B.4 Review Rubrics
@@
B.5 Reasoning Process Quality
@@
  • 2 (Poor): Some reasoning attempt but with major factual or procedural flaws; inconsistent or unclear Chain-of-Thought.
@@
B.6 Final Answer Accuracy
@@
Appendix C Model Detail Information
We provide the details of the evaluated models in Table 4.
Model Organization Size Backbone Source
Frontier Proprietary LLMs
GPT-5 OpenAI N/A gpt-5-2025-08-07
GPT-4.1 OpenAI N/A gpt-4.1-2025-04-14
GPT-5 mini OpenAI N/A gpt-5-mini-2025-08-07
GPT-4.1 mini OpenAI N/A gpt-4.1-mini-2025-04-14
Claude Sonnet 4.5 Anthropic N/A claude-sonnet-4-5-20250929
Claude Sonnet 4 Anthropic N/A claude-sonnet-4-20250514
Claude Sonnet 3.7 Anthropic N/A claude-3-7-sonnet-20250219
Gemini 2.5 Pro Google N/A Last Update: June 2025
Gemini 2.5 Flash Google N/A Last Update: June 2025
DeepSeek V3.2 DeepSeek N/A Last Update: Sep 29 2025
DeepSeek V3.1 DeepSeek N/A Last Update: Sep 22 2025
DeepSeek R1 DeepSeek N/A Last Update: Jan 20 2025
Grok 4 Heavy xAI N/A grok-4-0709
Grok 4 Fast xAI N/A grok-4-fast-reasoning
Finance Specific LLMs
Fino1 TheFinAI 8B meta-llama/Llama-3.1-8B TheFinAI/Fin-o1-8B
Fin-R1 SUFE-AIFLM-Lab 7B Qwen/Qwen2.5-7B-Instruct SUFE-AIFLM-Lab/Fin-R1
DianJin-R1 Qwen DianJin Team 7B Qwen/Qwen2.5-7B-Instruct DianJin/DianJin-R1-7B
Finance-LLaMA Wiro AI 8B deepseek-ai/DeepSeek-R1-Distill-Llama-8B WiroAI/WiroAI-Finance-Llama-8B
Finance-Qwen Wiro AI 7B Qwen/Qwen2.5-7B WiroAI/WiroAI-Finance-Qwen-7B
Math Enhanced LLMs
WizardMath WizardLM Team 7B mistralai/Mistral-7B-v0.1 WizardLMTeam/WizardMath-7B-V1.1
MetaMath MetaMath Project 7B EleutherAI/llemma 7b meta-math/MetaMath-7B-V1.0
Mathstral Mistral AI 7B mistralai/Mistral-7B-v0.1 mistralai/Mathstral-7B-v0.1
Qwen2.5-Math Qwen Team 7B Qwen/Qwen2.5-7B Qwen/Qwen2.5-Math-7B-Instruct
General Purpose Open LLMs
LLaMA 3.1 Meta 8B - meta-llama/Llama-3.1-8B
Qwen 2.5 Qwen Team 7B - Qwen/Qwen2.5-7B-Instruct
Qwen 3 Qwen Team 8B - Qwen/Qwen3-8B
Table 4: Details of the organization and model source (i.e. model version for proprietary models, and HuggingFace model name for open-source models) for the LLMs evaluated in FinChain.
Appendix D Metric Evaluation and Ablations
To better understand the behavior of our proposed metric, we conducted a series of ablation experiments and comparative analyses. All quantitative results reported in this section are benchmarked against expert human evaluations of reasoning quality (see § B.1 for expert details).
D.1 Ablations of the DTW-Based Metric
@@
To also capture the correctness of the final result, we also tried a weighted sum of the DTW-based scores and the final answer’s correctness. We used all possible $\alpha$’s in range 0.10.90.1-0.9 to identify the best proportion. Based on these experiments, the $\alpha$ of 0.1 resulted in the best Spearman $\rho$ and it was used for further comparison.
D.2 Comparative Evaluation
We also evaluated a range of traditional text-similarity and reasoning metrics, including ROUGE-2, ROUGE-L, step-level precision and recall (marked as ‘non-DTW’ in the table), BERTScore, and our weighted sum of DTWNormGate and final answer’s correctness (DTWNormGate+FAC in the table). Each metric was correlated with expert-assigned Reasoning Process Quality score. Table 5 summarizes the top Spearman correlations with expert process judgments.
@@ D.2 Comparative Evaluation
BERTScore 0.287
Table 5: Spearman correlation with expert evaluation of the reasoning process.
As additional validation, we also measure other correlation metrics: Kendall tau (which measures similarity of rankings) and Pearson correlation. As shown in Table 6 and Table 7, our suggested metric still holds high correlation position even with other correlation metrics.
Metric Pearson $\rho$
DTWNormGate+FAC 0.584
DTW F1 (Soft) 0.584
DTW Precision (Soft) 0.578
DTW F1 (Gate) 0.568
DTW Avg. Path Score (Gate) 0.556
DTW Recall (Gate) 0.550
Step Recall (non-DTW) 0.549
DTW Precision (Gate) 0.548
Step Precision (non-DTW) 0.530
DTWNormSoft 0.515
DTW Avg. Path Score (Soft) 0.515
DTW Recall (Soft) 0.511
DTWNormGate 0.492
ROUGE-L 0.421
ROUGE-2 0.420
BERTScore 0.300
Table 6: Pearson correlation with expert evaluation of the reasoning process.
Metric Kendall $\tau$
DTW Precision (Gate) 0.511
DTWNormGate+FAC 0.509
Step Precision (non-DTW) 0.505
DTW F1 (Gate) 0.503
DTWNormGate 0.503
DTW Precision (Soft) 0.494
DTW F1 (Soft) 0.487
Step Recall (non-DTW) 0.477
DTW Recall (Gate) 0.463
DTWNormSoft 0.460
DTW Avg. Path Score (Gate) 0.420
DTW Avg. Path Score (Soft) 0.401
DTW Recall (Soft) 0.389
ROUGE-2 0.360
ROUGE-L 0.330
BERTScore 0.213
Table 7: Kendall $\tau$ correlation with expert evaluation of the reasoning process.
D.3 Discussion
The DTW-based variants consistently achieve the highest correlation with expert judgments, with the Normalized DTW Alignment Score (Gate Mode) with FAC emerging as the most reliable indicator of reasoning faithfulness. Non-FAC and the “Soft” variant yields slightly lower but still strong correlations, suggesting that the gating formulation better captures strict consistency, while the soft variant provides smoother sensitivity to near-correct reasoning. Compared to traditional metrics such as ROUGE or simple step-level precision and recall, DTW captures not only semantic similarity but also the structural coherence and numerical consistency of reasoning chains. These results highlight the usefulness of our proposed metric.
D.4 Suggestions For Reproducibility
@@
The way in which the reasoning steps are parsed from the model’s output plays an important role in the quality estimation. In this work we combine both LLM-based parsing and parsing using regular expressions. We try to split the response on reasoning steps using regular expressions, which capture ‘Step X’-like patterns in the response, if such patterns are not found, we instruct LLM to split the answer on steps, copy-pasting the whole step string. We also use LLM to parse the final answer from the response. We acknowledge that such approaches depend on the prompting strategy, and there is a chance that other parsing and comparison methods (like the use of executable symbolic engine such as SymPy) can produce varying results.
For better reproducibility we share our parsing model details and prompts used to extract final answer and reasoning steps. We used GPT-4.1 model. Results were parsed using OpenAI’s output schema, which allowed us to avoid inconsistency in outputs. As a system prompt we used the following text:
You are a strict financial reasoning parser.
Your task is to convert a noisy, long,
conversational Chain-of-Thought solution
into a clean and structured JSON object.
Follow these rules exactly:
1. No Calculations
Do NOT compute anything.
Do NOT recompute numbers from the text.
Do NOT round, simplify,
or adjust any numeric value.
2. Number Extraction (CRITICAL)
You must copy numbers
EXACTLY as they appear in the text.
Keep the sign intact
(do NOT remove or change negative signs).
Keep decimal precision exactly as written.
Remove commas, dollar signs,
and percent signs only as formatting,
not as value changes.
3. Step Extraction
Identify only the actual reasoning steps.
Ignore filler text: summaries,
meta-comments, confidence statements,
restatements, chatter.
Produce concise step descriptions
capturing the logic of each step.
For each reasoning step, output an object with:
"index": <step_number>,
"text": "<short description of the step>",
"value": <numeric result of that step or null>
4. How to Determine value
If the step contains a
computation result, copy the
final numeric result in that step.
Usually this is the last
number in the step (e.g., number after "=").
If the step has no numeric result,
set "value": null.
5. Final Answer Extraction
"final_answer" must be copied exactly
as the last numeric value in the entire solution.
Apply the same numeric-copy rules as above.
6. JSON Output Format
Return ONLY a valid JSON object of the form:
"steps": [...],
"final_answer": <number>
7. Forbidden Behaviors
Do NOT change "-0.2636" into "2636.0".
Do NOT remove negative signs.
Do NOT round 1.514016 to 1.51.
Do NOT compute new numbers.
Do NOT invent numbers not present in the text.
Do NOT choose earlier numbers
if a later number is clearly the result.
Accuracy is measured strictly.
Copy numbers exactly.
As user prompt we used:
Here is the full solution text
to parse into reasoning steps
and a final answer:
[…]
Remember: follow the instructions and
return ONLY the JSON object.
In case if the steps were parsed with regular expression correctly, the final answer was retrieved using the following prompt:
You are given a full solution text.
Ignore step segmentation.
Your ONLY job is to extract
’final_answer’ as the last numeric
value in the solution,
following the same numeric-copy
rules from the system prompt.
Return ONLY a JSON object of the form:
"{ \"final_answer\": <number or null> }\"
Here is the solution text:
Appendix E Complementary Results
E.1 Results on a Sampled Subset
Due to the high inference cost of Grok 4 Heavy, we evaluate this model on a randomly sampled subset of 200 instances rather than the full benchmark. The results in Table 8 are reported to provide a coarse reference for this cost-limited setting and should not be directly compared with the full-benchmark results in Table 2.
Model ChainEval FAC ROUGE R2
Grok 4 Heavy 65.64 81.00 23.87
GPT-5 68.42 69.50 28.37
Gemini 2.5 Pro 67.92 73.00 18.85
Fin-R1 56.44 34.00 5.43
Mathstral 64.02 51.00 18.99
Table 8: Performance comparison on a randomly sampled 200-instance subset using ChainEval, final answer correctness (FAC), and ROUGE R2. These results are reported for reference only and are not directly comparable to full-benchmark results.
E.2 Complementary Domain-Level Results
We report additional domain-level results for the remaining models grouped by model category. These figures are provided for completeness and transparency, and no further analysis is conducted. Figure 8 reports domain-level ChainEval scores for additional frontier proprietary models, showing broadly consistent performance patterns across financial domains with moderate variation. Figure 9 presents domain-level performance for finance-tuned models, illustrating heterogeneous patterns across domains. Figure 10 shows domain-level results for math-enhanced models, with performance varying across financial domains. Figure 11 reports domain-level performance for general-purpose open models, serving as reference baselines across domains.
Figure 8: Domain-level performance of frontier proprietary models. Radar plot showing ChainEval scores across twelve financial domains for GPT-5, GPT-5-mini, Claude Sonnet 4.5, Gemini 2.5 Pro, and DeepSeek v3.2.
Figure 9: Domain-level performance of finance-tuned models. Radar plot showing ChainEval scores across twelve financial domains for Fin-R1, DianJin-R1, Fin-o1, Finance-LLaMA, and Finance-Qwen.
Figure 10: Domain-level performance of math-enhanced models. Radar plot showing ChainEval scores across twelve financial domains for Mathstral, Qwen2.5-Math, WizardMath, and MetaMath.
Figure 11: Domain-level performance of general-purpose open models. Radar plot showing ChainEval scores across twelve financial domains for LLaMA 3.1 Instruct, Qwen 2.5 Instruct, and Qwen 3.
E.3 Difficulty Breakdown under ChainEval
As shown in Figure 12, ChainEval exhibits a more gradual change across difficulty tiers than FAC. This indicates that models may maintain partial step-level alignment on harder instances even when end-task success decreases, consistent with ChainEval assigning partial credit to intermediate reasoning.
Figure 12: ChainEval across difficulty levels. ChainEval scores for representative models on Basic, Intermediate, and Advanced FinChain instances. Compared to FAC in the main text, ChainEval varies more gradually across difficulty, reflecting partial credit from step-level alignment.
Appendix F Error Analysis Supplementary
F.1 Error Taxonomy
This appendix defines the error categories used in the diagnostic error analysis. Each model output is assigned to a single category based on the primary criterion violated in the predicted reasoning or final result.
No Error. The reasoning process and final result match the gold computation within the predefined tolerance.
Computational. The applied formula or method is appropriate, but one or more intermediate or final numerical computations are incorrect.
Conceptual. The reasoning violates a problem constraint or applies a financial rule or formula inconsistently with the gold specification.
Sign / Direction. The sign or directional definition of a quantity is inconsistent with the gold formulation, such as reversing subtraction order or treating a decrease as an increase.
Unit / Precision. Units, scale, or numerical precision are handled inconsistently with the problem specification, for example percent versus decimal or dollars versus millions.
Parsing / Format. The output structure is malformed or inconsistent, preventing reliable parsing or alignment of reasoning steps.
Hallucination. One or more variables, assumptions, or quantities are introduced that are not supported by the original question or provided context.
F.2 Error Examples
Table 9 presents one representative example for each error category used in the diagnostic analysis. Each example consists of a short excerpt from the model output and a concise description of the specific criterion under which the output is labeled as erroneous. The descriptions focus on observable mismatches between the model output and the corresponding gold computation or definition, such as numerical inconsistency, violation of problem constraints, or unsupported quantities. The examples are intended solely to illustrate how the error taxonomy is applied in practice. They do not aim to explain the underlying causes of model behavior, assess error frequency, or attribute failures to model training, architecture, or reasoning capacity.
Error Type Model Model Output (snippet) Error Description
Computational GPT-5-mini “Total cash = 3,592,610.3 $\times$ 1.8632 = \$6,693,751.51.” Incorrect multiplication; correct total is approximately \$6.68M.
Conceptual GPT-5-mini “47408 + x = 0.8(60427 + x) $\Rightarrow$ x = 4668.” Violates conservation of total portfolio value during rebalancing.
Sign / Direction GPT-5-mini “Change = new - original $\approx$ -0.123.” Uses an inconsistent sign convention; gold definition is old - new.
Unit / Precision GPT-5-mini “175 (whole credits) … total rebate = 175 $\times$ 2.03 = \$355.25.” Applies unjustified rounding to a fractional quantity (175.5 credits).
Parsing / Format Finance-Qwen “… on the investment at the end of 3 years?” Output is malformed and cannot be reliably parsed into steps.
Hallucination Finance-Qwen “Assume 500M shares … Market cap = EPS $\times$ P/E = \$33.74B.” Introduces an unsupported quantity and applies an inconsistent valuation formula.
Table 9: Representative error examples with model outputs (expert-audited sample).